OpenStudy (anonymous):

An expression for the recommended weight of an adult is 11/2 (height in inches) - 220. These are my instructions. Now here's the problem: A woman weighs 110 pounds. What is her height in inches. I don't want to know the answer, I need somebody to set up the problem and explain it to me. Please help immediately, I need it for a test I am taking tomorrow morning at 9a.m. Thank You!(:

OpenStudy (anonymous):

you know that \[\frac{11}{2} \text{height} -220 = \text { weight}\] and so you have \[\frac{11}{2}\text{height } -200 = 110\]

OpenStudy (anonymous):

sorry should have been \[\frac{11}{2}\text{height } -220 = 110\]

OpenStudy (anonymous):

now you have to solve for "height" do you know how to do it or do you need help solving?

OpenStudy (anonymous):

I need help, i am totally confused(:

OpenStudy (anonymous):

ok but first do you understand what i wrote? you were given the weight, and an equation to find the weight given the height, so i set the equation equal to the weight you were give (110) and now we will find the height

OpenStudy (anonymous):

I understand that part

OpenStudy (anonymous):

we want to get the height by itself on one side of the equal sign, so the answer will look like \[\text{height} = \text {number}\]

OpenStudy (anonymous):

Yeah, so what do we do with the fraction, turn it to a mixed number?

OpenStudy (anonymous):

i will write out the steps. \[\frac{11}{2}\text{height } -220 = 110\] add 220 to both sides \[\frac{11}{2}\text{height } -220+220 = 110+220\] \[\frac{11}{2}\text{height } = 330\] multiply both sides by 2 \[2\times \frac{11}{2}\text{ height} =2\times 330\] \[11\times \text{height} = 660\] divide both sides by 11 \[\frac{11}{11}\times \text{height} = \frac{660}{11}\] \[\text{height} = 60\]

OpenStudy (anonymous):

Thank you so much!!!!!

OpenStudy (anonymous):

hope it is clear yw

OpenStudy (anonymous):

It is(: